OpenStudy (anonymous):

Carlos solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x - 2y = 10 y = 5 - 2x x - 2(5 - 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5

myininaya (myininaya):

what happen to -10 any why is it 10?

myininaya (myininaya):

x - 2(5 - 2x) = 10 x + 10 + 4x = 10 <-bad distributing

myininaya (myininaya):

so it was suppose to be x-10+4x=10 solve from there to get the right answer

OpenStudy (anonymous):

he miss the 10